UT 06-Feb-2020 D/E Night

ISA: B. Rothberg
  • DONE Complete LUCI-2/AGW-2 On-Sky Checks - high priority - J. Hill/B. Rothberg - 1-2 hours
    • See UT-2020-01-30 Wiki page for gory details
    • Files located on:
      • RM507-5/4: /home/observer/share/Service_Observing/Eng_Jan_2020/LUCI2_AGW_Checkout
      • OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/LUCI2_AGW_Checkout
    • DONE Verify daytime cable fix did not incur other issues - check off-axis guiding at various distances in the patrol field
    • DONE Verify S-H spots are centered - adjust if needed
    • DONE Measure Center of Rotation - adjust if needed
    • LUCI Detector Rotator Center (2 opposing PAs, 2 pairs if time)
    • DONE Verify Rotator ZP with Astrometric Masks - tolerates up to 1.2" seeing
      • INAF mask had 3 stars - only 1 was properly centered, updated mask was done wrong and did not work on-sky (PI used wrong version of LMS)
      • Rotator offset gave two different answers depending on machine used
        • RM507-5: Angle = -0.0479 err RA =0.073 err Dec = 0.123
        • OBS-4: Angle = -0.00072 Err RA = 0.073 and Err Dec = 0.134
        • Amended mask run by D. Thompson Angle = 0.09366, err RA=0.076 err Dec=0.051 - done offline UT 03-Feb-2020
        • Current Rotator ZP is 177.05 (this value was confirmed as the one used in 2019B - large offsets seemed to work for telluric)
    • DONE Take Confirmatory Transform Data - Stone Field - up to 1.5" seeing on DIMM
      • Send an Active preset to Stone D or Stone E, collimate
      • Then run transform_collect as in instructions
      • Config for LUCI-2
        • N30 FieldStop
        • 2.51sec DIT
        • H-band or FeII (take a test exposure in the field)
      • lists are located here on OBS-N: /home/lbto/idl/wfsc/transform/lists
      • use:
        • Stone_D_LUCI_PA0_short.list
        • Stone_Em135_0.list ( don't have to complete full list)
  • AO / IRTC Work - AO on-sky data collection tasks - pending release of IRTC (computer issues) and confirmation from AO group they need/want time - IN Parallel with LUCI-2/AGW-2 work above - D. Miller
  • Geo-Sync satellite - Anik F1R - requires IRTC on SX w/S
  • SOUL AO - LUCI-2 SL on DX - high priority - 2hours flexible
  • D-time Programs - depends on conditions and tasks completed - Moon at 81% illumination until 4am

OSA: J. Williams

ISp: G. Bechetti

Mtn Manager: P. Hartley

ISA: B. Rothberg (Tucson)

Techincal Wizard: J. Hill (Tucson)

Will hand over to PEPSI at 03:47 am local time

Startup

UT 00:20 - Init all on LUCI-2 - OK

UT 00:44 - still working on IRTC computer

L2 0057-0060 - 2.51 sec darks

UT 01:20 - IRTC is a no-go, could be fiber problem, serial cable, or power supply. Will have to wait until daytime to climb up in there and check.

UT 01:21 - we have a 10C differential on the mirrors - will open and check seeing, likely to go for bad seeing programs.

Going to backup plan C

UT 01:28 - going to pointing star and colliation - LUCI-2

-both M1s panicked, now only SX panicked

collimation star shows FWHM 3-4"

UT 01:50 - moving to Telluric for Arp 187

-aborting wait for collimation, i just need photons down the slit

-FWHM 3"

-offset +9.08" -1.01"

-error on committing offset 0, +10" - "guide loop didn't pause with timeout"

-manual offset failed 0,+10"

-restarting GCS

-resending preset

FWHM 2.6" on guider 1.8" on WFS

-skirting wind limits

-something block the bottom part of the 0.5" slit? - see it in both HKspec and zJspec maybe at -40" - weird since the top of the slit should be blocked due to guideprob

UT 02:24 - hung on 2nd offset (0,-10")

UT 02:32 - aborting waiti for preset, got through 1 exposure on the zJ spection

-restarting GCS

UT 02:34 - star not found - on edge - need pointing correction

UT 02:28 - resendng preset

UT 02:44 - skipping to the zJspec portion

L2 0064-0065 -telluric spectra HK

L2 0069-0072 -telluric spectra HK

L2 0073 -telluric spectra zJ

L2 0077-0080 - telluric spectra zJ

UT 02:51 - moving to Arp 187

-guideprobe motion taking a very long time....several minutes

-stuck? GCS hung again?, restarting GCS

UT 02:57 - resending preset

CONFIRMED the W-unit was NOT parked today - called Doug to park it.

UT 03:03 - resending preset since we lost guidestar

UT 03:14 - astigmatism also seems to have run away...Josh manually collimating

UT 03:31 - Putting mask in FPU

UT 03:33 - sending preset

UT 03:35 - resending - seems to only get every other preset

UT 03:53 - made a simple imaging script for field since we are testing GCS and want to avoid keeping mask in turnout

-note we're running on TCS2, could this be the problem?

UT 04:12 - reverting to previous TCS build

UT 04:25 - restored settings after reversion of TCS build - forgot to restart OAC and set shell

-resending preset

UT 04:28 - going to pointing star

-on edge of field out of focus, paddling to star

UT 04:36 - Closing for high winds - putting LUCI in safe mode

UT 04:43 - Light on WFS (AGW2) - 4Dlaser? -moon?

UT 07:02 - completed MODS calibrations - LUCI depends on if we re-open and what we do next

UT 07:11 - glass is -10.3C ambient is -12.5C - just need winds to die down

UT 07:34 - opening enclosure

AGW2 Checkout Attempt

UT 07:40 - oops, forgot to move the LUCI-2 cal unit back out of the way...doing it now

UT 07:50 - round in guider, FWHM 1.67"

-SH spots seem centered

-Attempt to measure Rotator Center

-need to move to a fainter star

Current Hotspot is X=270.0 Y=238.0

right_guider000002.fits - no star in image

UT 08:07 - resending active preset

right_guider000004.fits

On DS9: X=130 Y=113

multiply by 2 for binning: X=260 Y=226

Correct for DETSEC +15 X +39 Y --> X=275 Y=265

delta X= +5 Delta Y= 27
  • convert the desired correction from pixels to arcsec to mm (1 pixel = 1/20 arcsec for Potsdam AGw; 0.600 mm per arcsec)
Y=+0.81mm, this adjusts to RAD axis 0.81*1000 = +810 microns

need to adjust in oacontrol.conf:
AXIS_RAD
oofset = 418372 (old value) + 810, new value = 419182

-added notes

rdwrconfig -u [#]

UT 08:23 - Restart GCSRight, resending active preset

Verifying adjustment

right_guider000005.fits

X=129, Y=98 --> multiply by 2 for binning: X=258, Y=196 Correct for DETSEC +15 X +39 Y --> X=273, Y=235 (X=270, Y=238)

DONE

UT 08:32 - moving to Stone I for Rotator ZP

angle =0.10 but yellow in RA Star #35 has largest uncertainty in RA - removing it (RA err 0.101 Dec err 0.059)

angle = 0.09 (RA err 0.080, Dec err 0.051)

UT 09:04 - decision is to keep rotator ZP that is in place.

NOTE: THIS MOS ALIGNMENT WAS DONE ACROSS THE STREET DURING A FIRE ALARM ON MY LAPTOP

UT 09:07 - sending active preset to Stone I

UT 09:10 - putting N30Fieldstop into FPU

UT 09:18 - as a note for the Clamops, there were 4 x50 step corrections, then 2x100 step corrections

Test exposure to get enough flux in ring

H-band 2.51sec

will use Stone_I_m93.list for transform

L2 0121 - H-band 2.51sec DIT x2 NDIT (use this as dark) - holes are good

L2 0122 - moved guideprobe off axis will check star brightess

UT 09:22 - starting transform (DIMM now shows 2")

not doing ndits, so DIT=2.51 NDIT=1

-Note - seeing is all over the place

L2 0123 - GS 0

L2 0124 - dark

L2 0125 - GS 10

L2 0126 - GS 13

L2 0127 - GS 18

L2 0128 - GS 11

L2 0129 - GS 15

L2 0130 - GS 16

L2 0131 - GS 6

L2 0132 - GS 4

L2 0133 - GS 1 (guide-probe shadow?)

L2 0134 - GS 5

L2 0135 - GS 12

L2 0136 - GS 17

Failed to move to GS20

Backup Science Program - LUCI-2

Seeing is highly variable - NGC 3656 Local Merger Program

UT 09:49 - moving to Telluric star for NGC 3656

L2 0138-0141 airmass 1.042

UT 10:00 - moving to NGC 3656

L2 0142 -0148

seeing all over the place, ranging from 0.8 - 2.7" on guider and 1" to 2.5" on DIMM

-quicklook, I can see the CO bandhead, first two overtones

UT 10:42 - last exposure completed (got 30minutes) - LUCI-2 in safe config - handing over to PEPSI

UT 11:08 - MODS put to sleep

MODS calibrations

Slitless Pixel Flats

MODS1Blue taking a long time to write out exposure

M1B 0002-0011

M1R 0002-0006

M2B 0002-0011

M2R 0002-0006
Dual Grating 0.6" flat

M1B 0012-0017

M1R 0007-0012

M2B 0012-0017

M2R 0007-0012
Dual Grating Arcs

M1B 0018-0020

M1R 0013-0015 - 0015 seems saturated/odd - copied over from mods1data and also looks bad -- will have to re-take Updated IT 5862 - this might be my fault, i accidentally aborted the script and continued from the Kr+Xe, but the red exposure time is set earlier so it may have picked up the exposure from

--just going to ake the MODS-1 arcs again to be sure

M1B 0027-0029 (2nd set of arcs - use these)

M1R 0022-0024 (2nd set of arcs - use these)

M2B 0018-0020 (had to run blue fitsflush on last exposure)

M2R 0013-0015
Dual Grating 1.0" flat

M1B 0021-0026

M1R 0016-0021

M2B 0021-0026

M2R 0016-0021
Dual Grating 0.8" flats

M1B 0030-0035

M1R 0025-0029

M2B 0027-0032

M2R 0022-0027
8Kx3K Bias

M1B 0036-0040

M1R 0030-0034

M2B 0033-0038

M2R 0028-0032

LUCI Calibrations

zJspec G200 1" slit (O. Kuhn)

L2 0083-0087 zJ flats off

L2 0088-0092 -zJ flats on

L2 0093-0094 - Neon Arc

L2 0095-0096 - Argon Arc

HKSpec G200 1" slit (B. Rothberg)

L2 0097-0101 HK flats off

L2 0102-0106 HK lights on

L2 0107-0108 - Arcs lights off

L2 0109-0110 - Argon

L2 0111-0112 - Xenon

L2 0113-0114 - Neon

-- BarryRothberg - 05 Feb 2020
Topic revision: r4 - 06 Feb 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback